# Only Send What You Need: Learning to Communicate Efficiently in Federated Multilingual Machine Translation

### v1

@@ Abstract
Federated learning (FL) is a promising approach for solving multilingual tasks, potentially enabling clients with their own language-specific data to collaboratively construct a high-quality neural machine translation (NMT) model. However, communication constraints in practical network systems present challenges for exchanging large-scale NMT engines between FL parties. In this paper, we propose a meta-learning-based adaptive parameter selection methodology, MetaSend, that improves the communication efficiency of model transmissions from clients during FL-based multilingual NMT training. Our approach learns a dynamic threshold for filtering parameters prior to transmission without compromising the NMT model quality, based on the tensor deviations of clients between different FL rounds. Through experiments on two NMT datasets with different language distributions, we demonstrate that MetaSend obtains substantial improvements over baselines in translation quality in the presence of a limited communication budget.
1 Introduction
@@
Federated Learning (FL) has emerged as a popular distributed machine learning paradigm. FL enables collaborative model training among a set of clients via periodic aggregations of local models by a server McMahan et al. (2017); Konecný et al. (2016). The FL property of keeping data local to clients has important privacy advantages that have made it attractive for many learning applications.
@@
Natural language processing (NLP) is one domain standing to benefit from FL since user-generated text may contain sensitive information. Among the applications of FL in NLP, relatively few works have considered multilingual NLP and the impact of different languages on FL Liu et al. (2021). In recent years, neural machine translation (NMT) has shown substantial progress in this domain with the advent of large-scale language models such as BERT Devlin et al. (2019), GPT Radford et al. (2019), and their extensions. NMT has a further natural alignment with FL given its setting of non-IID local data distributions Weller et al. (2022): each client (user) typically has a specific language direction they are interested in for translation, which their local dataset will be skewed towards, motivating them to collaborate with each other via FL to construct the general NMT model.
@@
On the other hand, resource utilization is often a concern in deploying large-scale NMT models due to demands imposed on computational and memory resources Ganesh et al. (2020); Gupta and Agrawal (2020). While FL will distribute the processing load, every client must exchange its model parameters with a central server during the FL communication phases. Communication efficiency is a known bottleneck in traditional FL applications McMahan et al. (2017) and becomes an even more critical challenge with large-scale NMT models.
@@
In this paper, we are interested in optimizing multilingual NMT performance over an FL system with a limited communication budget. A premise for our work is that exchanging complete NMT engines in FL might not be necessary, similar to the argument in Passban et al. (2022). We can develop some intuition around this through a small FL experiment using the well-known FedAVG algorithm McMahan et al. (2017). In Figure 1, we perform FL on the UN Corpus dataset (see Section 4 for details) distributed across three clients (each containing one language translation direction), and plot the differences in NMT model tensors between a few consecutive training rounds for one of the clients. These differences are computed and visualized tensor-by-tensor, indicating the deviation of each tensor. We observe that the majority of deviation in the NMT model tensors cluster within small norms, while a small subset of tensors exhibit significant deviations. This observation is consistent across all clients, datasets, and data distribution combinations considered in this study.
@@
Recently, Passban et al. (2022) proposed a strategy of focusing on either highly fluctuating or less active tensors during FL communication to reduce computation load. Their approach involves sending only a fixed portion of NMT parameters – namely, by sending either the top 50% or the bottom 50% of tensors based on their deviation, which is computed using the previous round’s engine. The result of this is illustrated by the red threshold for the cases in Figure 1. However, sending a fixed portion of the parameters does not account for the fact that the deviation distribution will likely vary dynamically across rounds (as also observed in Figure 1). As a result, this approach is not sensitive to NMT quality, potentially resulting in the transmission of either too many parameters, i.e., extra communication burden without any significant change in translation quality, or too few parameters, i.e., leading to an undesirable model that negatively impacts translation quality.
@@
Contributions. Motivated by this, our objective is to explore dynamic thresholding techniques for multilingual federated NMT. The central challenge involved is how to adaptively determine a threshold that selectively filters out parameters from transmission until we expect that the translation quality will start to be compromised. To address this, our methodology, MetaSend, incorporates a meta-learning approach that generates a dynamic sending threshold adapting to the varying deviation distribution across training rounds. The result is depicted by the blue thresholds for the cases in Figure 1. In doing so, MetaSend considers translation quality and communication efficiency as important objectives in multilingual NMT training. In developing MetaSend, we make three major contributions:
@@
2 Related Work
2.1 Efficient NLP
@@
Previous research has explored efficiency enhancements for large NLP models from a computational perspective, i.e., achieving comparable results with fewer resources Treviso et al. (2022). Some studies have focused on the data side, e.g., showing how smart downsampling of available datasets can result in equal or improved performance compared to using the entire dataset Lee et al. (2021); Zhang et al. (2022). On the other hand, efforts to enhance efficiency through model designs include questioning the necessity of full attention heads in large language models and demonstrating that removing certain attention heads does not significantly impact test performance Kovaleva et al. (2019); Tay et al. (2020); Raganato et al. (2020). Compared to these works, motivated by the recent demand for FL in NLP, we specifically focus on communication efficiency in federated multilingual NMT and design a strategy that selectively transmits only the essential parameters of NMT engines for learning.
2.2 FL in NLP
@@
Recent research has begun exploring FL methods for NLP applications requiring privacy preservation Sui et al. (2021); Qin et al. (2021); Ge et al. (2020); Lin et al. (2022). During the FL communication phase, large NLP models are exchanged, introducing a significant communication cost associated with model updates. To address this, Melas-Kyriazi and Wang (2022) proposed a gradient compression methodology for language modeling, while Ro et al. (2022) leveraged similarities between smaller and larger models in cross-device FL. The parameter selection method proposed in Passban et al. (2022) targeting mixed-domain NMT is most relevant to our work. As discussed in Section 1, they analyze how much deviation each client has from its previous round and send a fixed-portion of model’s tensors during FL communication. However, sending a fixed amount of tensors without considering the varying deviation distribution can lead to lower translation quality by potentially removing meaningful parameters. To handle these challenges, we design a meta-learning-based parameter selection method that sends the client’s parameters with dynamic thresholds to capture the tradeoff between communication efficiency and translation quality.
3 MetaSend for Federated NMT
3.1 Problem Setup: Federated NMT
@@
When training the NMT model over multiple clients, we follow the general cross-silo FL setting introduced by Li et al. (2021). Algorithm 1 summarizes the overall training procedure. Each client first runs stochastic gradient descent (SGD) on its local data and then Sends the learned NMT model to the server. The server then executes a global aggregation after receiving all the trained models. We assume the standard FedAVG algorithm McMahan et al. (2017) as the aggregation method. Using FedAVG as opposed to other aggregation procedures allow us to focus on the communication efficiency of NMT in the FL scenario.
@@
3.2 Overview of MetaSend
During FL communication, a large amount of NMT model weights have to be uploaded to the server during the Send action (in Algorithm 1) for aggregation. This communication can be quite costly for a large NMT model, which is a key bottleneck for FL. To tackle this challenge, we propose MetaSend, which adapts the NMT tensors sent based on a customized sending threshold for each communication round. The key idea of MetaSend is to build Model-Agnostic Meta-Learning (MAML) Shu et al. (2019) into the FL rounds to balance communication efficiency and translation quality.
@@ 3.2 Overview of MetaSend
3.3 MetaSend: Customized Sending and Aggregation
In this subsection, we answer the first question mentioned above. Our intuition is that the extent of model parameter deviation relative to the original norm provides an indication of whether information is worth sending. Our observation in Section 1 shows that the tensors of the NMT model responsible for learning exhibit a clustered pattern in the deviation distribution.
@@ 3.3 MetaSend: Customized Sending and Aggregation
3.4 MetaSend: MAML Module Update
@@
To address the second question, we aim for our MAML module to generate an adaptive sending threshold based on translation quality. Figure 3 shows the learning process and the optimization flow of this module. represents the hyperparameter set of our MAML module in communication round , and is the generated threshold from the module. After the parameter selection process and aggregation (Equations 3 and 4), the parameters of the -th resulting client model and the global model can be expressed as and , respectively. To assess the quality of the global model , we randomly select batches of samples from the validation dataset and evaluate the global model using these samples. Subsequently, we employ the validation loss as the optimization objective for the MAML module, which encourages the module to adapt in the direction of superior translation quality. Thus, our MAML module update can be written as:
@@
where is the learning rate for the meta update. By optimizing the MAML module with consideration of the translation quality of the global NMT model, our MAML module can generate a customized threshold for each round that considers both the deviation distribution and the translation quality. We will see in Section 5 how this process of learning what parameters to send results in substantial translation quality and communication efficiency improvements.
4 Experimental Design
4.1 Datasets and Client Partitioning
We utilize two NMT datasets: MTNT Michel and Neubig (2018) and UN Corpus Ziemski et al. (2016). The MTNT dataset comprises English to French (En Fr) and English to Japanese (En Ja) translations, while the UNMT dataset includes three official UN language directions: English to French (En Fr), Arabic to Spanish (Ar Es), and Russian to Chinese (Ru Zh). Three training settings are considered: (i) centralized training without FL, (ii) FL with IID data, where the data for each client is sampled randomly from all data, and (iii) FL with non-IID data, where each client only sees data for one language direction. See Appendix A for more details on the datasets.
4.2 Base Model and Evaluation Metrics
Following the multilingual FL experimental settings in Weller et al. (2022), we use the M2M-100 model Fan et al. (2021) to conduct machine translation. The M2M-100 model is a sequence-to-sequence model with 418M parameters, and it can translate between any pair of 100 languages. We measure the quality of translation by using sacreBLEU Post (2018) and COMET Rei et al. (2020). SacreBLEU is a commonly used metric for evaluating NMT quality, while COMET is a more advanced metric that shows some degree of correlation with human judgement. See Appendix A for more details on these translation metrics.
Besides translation quality, we consider two metrics for FL efficiency: tensor saving and processing time. Tensor saving is defined as the ratio of tensors that are not exchanged between the server and clients during the Sending step in Algorithm 1 (or line 7 in Algorithm 2). For efficiency evaluation, we will report the average tensor savings and the exact processing time over all training rounds.
@@ 4.2 Base Model and Evaluation Metrics
| 31.1 | 14.6 | 22.9 | 37.9 | 38.5 | 38.7 | 38.4 | ||
| 31.9 | 15.5 | 23.7 | 38.4 | 38.7 | 39.2 | 38.8 | ||
| MetaSend (ours) | 32.0 | 16.8 | 24.4 | 42.0 | 39.1 | 41.9 | 41.0 | |
| MetaSend (ours) | 32.7 | 17.3 | 25.0 | 42.3 | 39.3 | 41.6 | 41.1 | |
@@
| 19.0 | 12.6 | 15.8 | 29.8 | 28.3 | 26.5 | 28.2 | ||
| 18.3 | 13.1 | 15.7 | 30.5 | 29.9 | 26.1 | 28.8 | ||
| MetaSend (ours) | 19.7 | 13.9 | 16.8 | 31.2 | 30.1 | 29.7 | 30.3 | |
| MetaSend (ours) | 20.1 | 13.6 | 16.9 | 32.8 | 31.3 | 29.7 | 31.3 | |
4.3 FL Training and MAML Module

### v2

@@ Abstract
Federated learning (FL) is a promising distributed machine learning paradigm that enables multiple clients to collaboratively train a global model. In this paper, we focus on a practical federated multilingual learning setup where clients with their own language-specific data aim to collaboratively construct a high-quality neural machine translation (NMT) model. However, communication constraints in practical network systems present challenges for exchanging large-scale NMT engines between FL parties. We propose a meta-learning-based adaptive parameter selection methodology, MetaSend, that improves the communication efficiency of model transmissions from clients during FL-based multilingual NMT training. Our approach learns a dynamic threshold for filtering parameters prior to transmission without compromising the NMT model quality, based on the tensor deviations of clients between different FL rounds. Through experiments on two NMT datasets with different language distributions, we demonstrate that MetaSend obtains substantial improvements over baselines in translation quality in the presence of a limited communication budget.
Index Terms:
Federated Learning, Machine TranslationI Introduction
@@
Federated learning (FL) has recently emerged as a popular distributed machine learning paradigm as it enables collaborative model training among a set of clients via periodic aggregations of local models by a server [2, 3]. The standard operation of FL is that each client sends its model parameters to the server, which periodically aggregates these parameters into a global model and synchronizes the global model to all clients in each communication round. The FL property of keeping data local to clients has important privacy advantages that have made it attractive for many learning applications.
@@
Natural language processing (NLP) is one domain standing to benefit from FL since user-generated text may contain sensitive information. Among the applications of FL in NLP, relatively few prior works have considered multilingual NLP and the impact of different languages on FL [4] despite their practical significance. In recent years, neural machine translation (NMT) [5, 6, 7] has shown substantial progress in this domain with the advent of large-scale language models such as BERT [8], GPT [9], and their extensions. NMT has a further natural alignment with FL given its setting of non-IID local data distributions [10]: each client (user) typically has a specific language direction they are interested in for translation, which their local dataset will be skewed towards, motivating them to collaborate with each other via FL to construct the general NMT model.
However, multilingual NMT presents unique challenges for FL compared with more generic learning tasks. The presence of language heterogeneity, due to variations in vocabulary, grammar structures, and syntax across clients and languages, complicates model training [10]. Dialects and regional variations add further complexity, as standard models may struggle to generalize across diverse linguistic patterns [4]. Dataset imbalance also manifests in unique ways for NMT tasks, with certain languages being underrepresented across clients and having significantly fewer resources devoted to them compared to others. Together, these factors can lead to performance disparities for different clients and languages.
@@
Additionally, resource utilization is often a concern in deploying large-scale NMT models due to demands imposed on computational and memory resources [11]. While FL will distribute the processing load, every client must exchange its model parameters with a central server during the FL communication phases. Communication efficiency is a known bottleneck in traditional FL applications [2] and becomes an even more critical challenge with large-scale NMT models.
I-A Communication Efficiency in Federated NMT
In this paper, we are interested in addressing the aforementioned challenges simultaneously, i.e., optimizing multilingual NMT performance over an FL system with a limited communication budget. While many existing works have studied communication-efficient FL for general learning tasks [12, 13], federated multilingual NMT is nuanced due to the language heterogeneity, dialects, and imbalance characteristics mentioned above. The recent work [14] was the first to explore this, arguing that exchanging complete NMT engines in FL may be unnecessary due to how deviations in parameters/tensors tend to be distributed across consecutive training rounds.
@@ I-A Communication Efficiency in Federated NMT
To build intuition around this, we conduct a small NMT FL experiment using the well-known FedAVG algorithm [2]. In Figure 1, we perform FL on the UN Corpus dataset (see Section IV for details) distributed across three clients (each containing one language translation direction), and plot the differences in NMT model tensors between a few consecutive training rounds for one of the clients. These differences are computed and visualized tensor-by-tensor, showing deviations for each tensor. We observe that most NMT model tensors have small deviations, while a small subset of tensors exhibit significant deviations. We find this to be the case across other clients, datasets, and distributions as well. This motivates parameter selection methods that send only specific tensors during each FL training round, with the selection tailored to this phenomenon of clustered deviations in federated NMT.
@@
The method in [14] aims to account for this by focusing on transmitting either highly fluctuating or less active tensors during FL rounds to reduce communication load. Their approach involves sending a fixed portion of model parameters – namely, by sending either the top 50% or the bottom 50% of tensors based on their deviation, which is computed using the previous round’s NMT engine. However, it is important to note that sending a fixed portion of the parameters does not account for the fact that the deviation distribution will likely vary dynamically across rounds (as also observed in Figure 1). As a result, this approach does not pay careful attention to the tradeoff between communication efficiency and NMT quality, potentially resulting in the transmission of either too many parameters, i.e., extra communication burden without any significant change in translation quality, or too few parameters, i.e., leading to an undesirable model that negatively impacts translation quality. Motivated by this, we pose the following research question: How can we adaptively adjust the parameters transmitted across training rounds in federated NMT to reduce communication burden while not compromising the translation model quality?
@@
To address this question, we focus on developing a dynamic thresholding technique that can be applied for NMT model transmission during FL. The central challenge involved is how to adaptively determine a threshold that selectively filters out parameters from transmission until we expect that the translation quality will start to be compromised. Moreover, our approach aims to address the sensitivity of NMT quality and mitigate the risk of either excessive communication burden or inadequate parameter exchange. To this end, we propose MetaSend, a model-agnostic meta-learning (MAML)-based methodology that generates a dynamic transmission threshold adapting to the varying tensor deviation distributions across training rounds. By crafting the MAML module adaptation according to the observed NMT quality across clients, our method accounts for the tradeoff between translation quality and communication efficiency in parameter/tensor transmission. The impact of this can be seen by the blue thresholds in Figure 1. Compared to sending a fixed portion (red threshold) as in [14], which can lead to sending too much (i.e., excessive communication) or too little (i.e., compromised NMT quality), our dynamic threshold accounts for varying NMT quality.
I-B Summary of Contributions
Overall, we make the following major contributions in developing MetaSend:
@@ I-B Summary of Contributions
The rest of the paper is structured as follows. Section II reviews previous work on efficient methods and federated learning for NLP. Section III presents the problem setup of federated machine translation and introduces our proposed method, MetaSend, for enhancing efficiency. Section IV provides details about the experimental setups, while extensive experimental results are presented in Section V. Finally, we conclude our work in Section VI.
II Related Work
II-A Computational Efficiency in NLP
@@
Previous research has explored efficiency enhancements for large NLP models from a computational perspective, i.e., achieving comparable results with fewer resources [15]. Some studies have focused on the data side, e.g., showing how smart downsampling of available datasets can result in equal or improved performance compared to using the entire dataset [16, 17]. On the other hand, efforts to enhance efficiency through model designs include questioning the necessity of full attention heads in large language models and demonstrating that removing certain attention heads does not significantly impact test performance [18, 19]. Several previous works have also investigated model pruning methods to improve the efficiency of learning large-scale language models, which identify and remove non-essential portions of networks. Earlier work focused mainly on unstructured pruning, where weights are pruned individually [20, 21]. More recently, structured pruning removes groups of consecutive parameters or structured blocks of weights, leading to significant speedups [22, 23]. Compared to these works, motivated by the recent demand for FL in NLP, we focus on communication efficiency in federated multilingual NMT and design a strategy that selectively transmits only the essential parameters of NMT engines for learning.
II-B Communication Efficiency in FL
Communication efficiency is a critical aspect of FL, particularly in scenarios with limited bandwidth and resource-constrained devices (see [24] for a recent survey). Extensive research has focused on enhancing communication efficiency through model and gradient compression techniques [25, 26], alternative communication topologies [27, 28], and intelligent parameter selection and sparsification methods [25, 29]. Our work on selective parameter/tensor transmission is most relevant to the selection/sparsification category [14, 25, 29]. Specifically, [14, 25] sort the model parameters by absolute magnitude and select a subset of parameters that rank among the highest or lowest in the entire model, while [29] uses historical model parameters to determine the important parameters for communication in the current model. Different from these works, MetaSend is intended specifically for NMT tasks, which as discussed in Section I exhibit nuanced tensor deviations across training rounds due to language heterogeneity, regional variations, data imbalances, and other disparities across clients. Unlike these methods that consider only communication efficiency in their design, our method optimizes communication based on the model’s performance, taking both performance and efficiency into account during the FL process. We will see in Section V that our approach outperforms baselines [25, 29], intended for more generic FL tasks, in the NMT setting we focus on.
II-C Federated Learning and NLP
@@
Recent research has begun exploring FL methods for NLP applications requiring privacy preservation [30, 31, 32, 33]. During the FL communication phase, large NLP models are exchanged, introducing a significant communication cost associated with model updates. To address this, [34] proposed a gradient compression methodology for language modeling, while [35] leveraged similarities between smaller and larger models in cross-device FL. The parameter selection method proposed in [14] targeting mixed-domain NMT is most relevant to our work. As discussed in Section I, they analyze client deviations from the previous round and send a fixed portion of model tensors during FL communication. However, this fixed approach may overlook meaningful parameters, reducing translation quality. To address this, we design a meta-learning-based parameter selection method with dynamic thresholds to balance communication efficiency and translation quality.
III MetaSend for Federated NMT
III-A Problem Setup: Federated NMT
@@
When training the NMT model over multiple clients, we follow the general cross-silo FL setting introduced by [36]. Algorithm 1 summarizes the overall training procedure. Each client first runs stochastic gradient descent (SGD) on its local data and then Sends the learned NMT model to the server. The server then executes a global aggregation after receiving all the trained models. We assume the standard FedAVG algorithm [2] as the aggregation method. Using FedAVG as opposed to other aggregation procedures allow us to focus on the communication efficiency of NMT in the FL scenario.
@@
III-B Overview of MetaSend
During FL communication, a large amount of NMT model weights have to be uploaded to the server during the Send action (in Algorithm 1) for aggregation. This communication can be quite costly for a large NMT model, which is a key bottleneck for FL. To tackle this challenge, we propose MetaSend, which adapts the NMT tensors sent based on a customized sending threshold for each communication round. The key idea of MetaSend is to build Model-Agnostic Meta-Learning (MAML) [37] into the FL rounds to balance communication efficiency and translation quality.
@@ III-B Overview of MetaSend
III-C MetaSend: Customized Sending and Aggregation
In this subsection, we answer the first question mentioned above. Our intuition is that the extent of parameter deviation relative to the original norm provides an indication of whether information is worth sending. Our observation in Section I shows that the tensors of the NMT model responsible for learning exhibit a clustered pattern in the deviation distribution.
@@ III-C MetaSend: Customized Sending and Aggregation
III-D MetaSend: MAML Module Update
@@
To address the second question, our MAML module generates an adaptive sending threshold based on translation quality, as illustrated in Figure 3. represents the hyperparameter set of our MAML module in communication round , and is the generated threshold from the module. After the parameter selection process and aggregation (Equations 3 and 4), the parameters of the -th resulting client model and the global model can be expressed as and , respectively. To assess the quality of the global model , we randomly select batches of samples from the validation dataset and evaluate the global model using these samples. Subsequently, we employ the validation loss as the MAML module’s optimization objective to guide it toward improving translation quality. Thus, our MAML module update can be written as:
@@
where is the learning rate for the meta update. It is important to note that although we are evaluating the global model, the validation loss is actually calculated using the validation dataset present in each client. This process can be seen as occurring after broadcasting the global model to each client at every round. In this process, each client computes the validation loss using the global model and sends the validation loss back to the server. Subsequently, the server updates the MAML module according to the aggregated loss. Finally, by optimizing the MAML module with consideration of the translation quality of the global NMT model, our MAML module can generate a customized threshold for each round that considers both the deviation distribution and the translation quality. We will see in Section V how this process of learning what parameters to send results in substantial translation quality and communication efficiency improvements.
IV Experimental Design
IV-A Datasets and Client Partitioning
We utilize two widely used NMT datasets: Machine Translation of Noisy (MTNT) dataset [38] and UN Machine Translation (UNMT) Corpus [39]. The MTNT dataset was gathered from user comments on Reddit. The dataset contains two language directions: English to French (En Fr) and English to Japanese (En Ja). The dataset contains 5,605 instances in each direction for training and 1k each for validation and test set. MTNT features diverse language pairs, each with unique linguistic characteristics, making it a strong candidate for testing the ability of our methodology to handle various grammatical, syntactic, and semantic challenges. The UNMT Corpus consists of manually translated UN documents, and we consider three language directions: English to French (En Fr), Arabic to Spanish (Ar Es), and Russian to Chinese (Ru Zh). The dataset contains 80k instances in each direction for training and 10k each for validation and test set. Similar to MTNT, UNMT captures a wide range of linguistic variations across its six languages, each with unique grammar, syntax, and semantic structures. For each dataset, we consider three training settings: (i) centralized training without FL, (ii) FL with IID data, where the data for each client is sampled randomly from all data, and (iii) FL with non-IID data, where each client only sees data for one language direction.
@@ IV-A Datasets and Client Partitioning
| DPg | 31.1 | 14.6 | 22.9 | 37.9 | 38.5 | 38.7 | 38.4 | |
| DPl | 31.9 | 15.5 | 23.7 | 38.4 | 38.7 | 39.2 | 38.8 | |
| MetaSendg (ours) | 32.0 | 16.8 | 24.4 | 42.0 | 39.1 | 41.9 | 41.0 | |
| MetaSendl (ours) | 32.7 | 17.3 | 25.0 | 42.3 | 39.3 | 41.6 | 41.1 | |
@@
| DPg | 19.0 | 12.6 | 15.8 | 29.8 | 28.3 | 26.5 | 28.2 | |
| DPl | 18.3 | 13.1 | 15.7 | 30.5 | 29.9 | 26.1 | 28.8 | |
| MetaSendg (ours) | 19.7 | 13.9 | 16.8 | 31.2 | 30.1 | 29.7 | 30.3 | |
| MetaSendl (ours) | 20.1 | 13.6 | 16.9 | 32.8 | 31.3 | 29.7 | 31.3 | |
IV-B Base Model and Evaluation Metrics
Following the multilingual FL experimental settings in [10], we use the M2M-100 model [40] to conduct machine translation. The M2M-100 model is a sequence-to-sequence model with 418M parameters, and it can translate between any pair of 100 languages. We measure the quality of translation by using sacreBLEU [41] and COMET [42]. SacreBLEU is a commonly used metric for evaluating NMT quality, while several research works have mentioned that the traditional overlap-based evaluation metrics do not correlate well with human evaluation [43, 44, 45]. Therefore, we also evaluated each method using a more advanced MT evaluation metric, COMET, demonstrating that the evaluated quality is better aligned with human judgement. Besides translation quality, we consider two metrics for FL efficiency: tensor saving and processing time. Tensor saving is defined as the ratio of tensors that are not exchanged between the server and clients during the Sending step in Algorithm 1 (or line 7 in Algorithm 2). For efficiency evaluation, we will report the average tensor savings and the exact processing time over all training rounds.
IV-C FL Training and MAML Module